Contribution Margin Harry Company sells 31,000 units at $22 per unit. Variable costs are $12.76 per unit, and fixed costs are $131,800 Determine (a) the contribution margin ratio, (b) the unit contribution margin, and (c) income from operations. a. Contribution margin ratio (Enter as a whole number.) b. Unit contribution margin (Round to the nearest cent.) c. Income from operations 0 per unit
Answer #1

a. Contribution Margin Ratio = (Contribution per unit / Selling price per unit)*100

Contribution per unit = Selling price per unit - Variable cost per unit

= 22 - 12.76

= $ 9.24 per unit

Hence, Contribution Margin Ratio = (9.24 / 22)*100

= 42%

b. Unit Contribution Margin per unit = Selling price per unit - Variable cost per unit

= 22-12.76

= $ 9.24 per unit

c. Income from Operations = Contribution - Fixed Cost

Contribution = Sales - Variable Cost

= (31000*22) - (31000*12.76)

= 682000 - 395560

= 286440

Therefore, Income From Operations =286440 - 131800

= $ 154640